OpenStudy (anonymous):

What is the midpoint of the line segment whose endpoints are (-8, 12) and (-13, -2)?

OpenStudy (anonymous):

\[(\frac{ -21 }{ 2 }, 5)\]

OpenStudy (anonymous):

the possible answers are (-10.5, 5) and (-10.5, 7)

OpenStudy (anonymous):

-8 - (-13) = 5, so would it be the first one?

OpenStudy (anonymous):

It's (-10.5,5)

OpenStudy (anonymous):

How did you get to that?

OpenStudy (anonymous):

I used the midpoint formula which is \[(\frac{ X1+X2 }{ 2 },\frac{ Y1+Y2 }{ 2 })\] I then plugged in the coordinates and solved it

OpenStudy (anonymous):

x1 is -8 and x2 is -13, right? @hlyweiman

OpenStudy (anonymous):

Yep. Makes sense!!!! Thank you sooooo much! :)

OpenStudy (anonymous):

Yep! You're more than welcome